OpenStudy (sdlrxo):

The molecular formula of glucose is C6H12O6. What would be the molecular formula for a polymer made by linking five glucose molecules together by dehydration reactions?

OpenStudy (madangerzone):

\(\ 6 * 10 = 60\) \(\ 10 * 10 = 100\) \(\ 5 * 10 = 50\) \(\ C60 ~ H100 ~O51 \) \(\ C60 ~H102 ~O51 \)